# Aligning MLLM Benchmark With Human Preferences via Structural Equation Modeling

### v1

@@ 2.3 A SEM-based Evaluation Framework for Benchmark
where $n$ denotes the total number of task indicators in the benchmark. This aggregate score $D_{\mathrm{valid}}\in(0,1]$ reflects the overall degree of indicator independence across the benchmark. Lower values indicate potential redundancy and multicollinearity issues at a global level, while higher values suggest stronger statistical validity of the task indicators as formative constructs.
3 Gold Benchmark
3.1 Hierarchical Structure Based on Piaget’s Developmental Theory
@@
We propose a new capability hierarchy grounded in Jean Piaget’s theory of cognitive development, a well-established framework in developmental psychology. Unlike previous heuristic-based groupings, the proposed approach introduces a human-centered, theory-driven structure to model the cognitive underpinnings of MLLM performance. By anchoring the evaluation framework in developmental principles, we aim to achieve a more interpretable, hierarchical, and cognitively faithful representation of model abilities.
@@
This Piaget-inspired cognitive structure is embedded into our benchmark design, with each task aligned to one of the three hierarchical cognitive layers (Perception, Memory, and Reasoning). SEM confirms the distinctiveness and validity of this stratification, demonstrating improved dimensional separability, reduced task redundancy, and stronger measurement consistency compared to conventional heuristic groupings.
3.2 Gold Construction
@@
Hyperparameters: VIF threshold $\delta_{\text{VIF}}$, loading threshold $\lambda_{\min}$
Step 1: Define latent variables: Perception, Memory, Reasoning
Step 2: Assign tasks $x_{i}\in\mathcal{X}$ to latent constructs
Step 3: Initialize weights $w\leftarrow 1$
while there exists task with VIF $>\delta_{\text{VIF}}$ or outer loading $<\lambda_{\min}$ do
     Estimate model via PLS-SEM (maximize explained variance)
     Remove tasks with VIF $>\delta_{\text{VIF}}$ (redundancy)
     Remove tasks with outer loading $<\lambda_{\min}$ (low quality)
end while
Step 3: Validate model with Cronbach’s $\alpha$ and HTMT
Output: A theory-aligned benchmark structure with validated task indicators
To construct a human-aligned benchmark structure, we map tasks from existing multimodal datasets to cognitive processes defined in a Piaget-inspired framework (Section 3.1). This process is formalized by using SEM, as outlined in Algorithm 1. We begin by curating an indicator pool by collecting tasks from four widely adopted multimodal benchmarks: MME fu2023mme , MMMU-Bench yue2024mmmu , BLINK fu2024blink and MMBench liu2024mmbench . Each task is manually annotated and assigned to one cognitive layer based on its processing demands (e.g., OCR $\rightarrow$ Perception; landmark recall $\rightarrow$ Memory; domain-specific inference $\rightarrow$ Reasoning).
Each task is manually annotated and aligned to one of the three cognitive domains based on its processing demands. We then apply Partial Least Squares SEM (PLS-SEM) to model the relationships among latent constructs and their associated indicators. This method is well-suited to our context, where sample size is limited and indicators may be collinear.
To refine the model, we iteratively prune the indicator pool by removing tasks exhibiting high multicollinearity (variance inflation factor $>\delta_{\text{VIF}}$) or insufficient indicator quality (outer loading $<\lambda_{\min}$). Following prior empirical standards fong2013hair ; 10990127 , we set $\delta_{\text{VIF}}=5$ and $\lambda_{\min}=0.75$. This process enhances both statistical validity and conceptual coherence. We further evaluate construct reliability and discriminant validity using Cronbach’s $\alpha$ and the heterotrait–monotrait (HTMT) ratio.
@@
3.3 Statistics.
Figure 4: Structural analysis of Gold by using SEM-based framework.
The final Gold Benchmark contains 12 tasks, selected from an initial pool of 85 across the four benchmarks. Our analysis follows the guidelines proposed by J. Hair hair2019use . For measurement model, we examine the indicator weights to evaluate their contribution to the latent constructs. As depicted in the figure, the weights are consistently greater than 0.75. This outcome indicates that the retained indicators possess strong explanatory power and are effectively capturing the underlying constructs they are intended to measure.
Table 1: Reliability and Validity Metrics for Latent Constructs
Construct Cronbach’s $\alpha$ Reliability Composite reliability Convergent validity
Memory 0.702 0.732 0.823 0.607
Perception 0.804 0.813 0.885 0.720
Reasoning 0.902 0.906 0.927 0.719
To further verify the statistical quality of the benchmark, we conduct standard reliability and validity tests on the latent constructs. As shown in Table 1, all constructs meet the recommended thresholds for internal consistency and convergent validity. Specifically, Cronbach’s $\alpha$ values are all above 0.70, reflecting acceptable internal reliability. Composite reliability scores exceed 0.80 across all dimensions, confirming the coherence of the grouped indicators. In terms of convergent validity, all constructs exhibit Average Variance Extracted (AVE) values above 0.60, suggesting that each construct explains a substantial portion of the variance in its associated indicators.
We also assess multicollinearity using the Variance Inflation Factor (VIF). As shown in Figure 4(b), all indicators yield VIF values well below the conservative threshold of 5 (ranging from 1.12 to 2.75), suggesting the absence of problematic multicollinearity and supporting the discriminant structure of the model. Finally, we evaluate discriminant validity using the Heterotrait-Monotrait (HTMT) ratio in Figure 4(c). All pairwise HTMT values are below the recommended cutoff of 0.90, with the highest observed value being 0.86 (between Memory and Perception). This result confirms that each latent construct captures a distinct cognitive domain and is not confounded with others.
These statistical results collectively validate the robustness of our benchmark construction. They demonstrate that the selected tasks reliably capture the latent cognitive dimensions defined in our Piaget-inspired framework, thus ensuring both empirical rigor and theoretical alignment.
4 Experimental Study
4.1 Experiment Setting
To evaluate the quality and interpretability of the Gold benchmark, we design a series of experiments involving both open-source and closed-source MLLMs. Our evaluation includes two core settings: (1) Main Experiment, where we assess mainstream models using a standardized pipeline to reveal overall performance differences; (2) SEM-based Benchmark Evaluation Experiment, where we validate Gold’s diagnostic effectiveness by measuring key metrics such as modality diversity ($D_{div}$), task contribution ($TC$), and diagnostic validity ($D_{valid}$). We further assess its interpretability by comparing these outputs with human ratings to examine alignment with human judgment.
Evaluation Models. For main experiment, we evaluate both closed-source and open-source MLLMs. For the closed-source group, we include GPT-4o achiam2023gpt . For the open-source group, we consider a range of models with varying scales, including Qwen2.5-VL bai2025qwen2 , Gemma3 team2025gemma , and InternVL3 zhu2025internvl3 . These models represent the current frontier of multimodal LLM development and cover diverse architecture families and parameter sizes. For the benchmark evaluation experiment, we leverage evaluation results from 165 multimodal language models collected via VLMEvalKit duan2024vlmevalkit , which provides standardized testing across multiple benchmarks. The Metrics used in the experiment will be described in detail in the Appendixes.
Table 2: Comparison of state-of-the-art methods on the Gold. BMK and CS denote Basic Medical Knowledge and Computer Science, respectively. Overall scores are computed by weighting capability-level scores according to their corresponding task contributions.
Methods Model size Overall Perception Memory Reasoning
@@ 4.1 Experiment Setting
[Uncaptioned image]  GPT-4o - 165.7 92.5 92.5 96.3 72.6 91.0 73.3 56.7 73.3 80.0 76.7 70.0
@@
[Uncaptioned image]  Gemma3 4B 118.6 54.2 55.0 81.3 62.5 80.1 53.3 36.7 46.7 46.7 53.3 50.0
[Uncaptioned image]  Gemma3 12B 139.4 75.0 75.0 85.0 75.6 88.3 63.3 33.3 60.0 73.3 63.3 53.3
[Uncaptioned image]  InternVL3 2B 116.9 86.7 77.5 85.0 78.1 84.5 30.0 40.0 40.0 56.7 50.0 40.0
[Uncaptioned image]  InternVL3 9B 153.6 89.2 86.7 96.3 80.9 87.9 63.3 66.7 63.3 60 73.3 56.7
[Uncaptioned image]  InternVL3 38B 163.7 86.7 82.5 96.3 82.4 92.3 76.7 50.0 70.0 76.7 83.3 63.3
[Uncaptioned image]  InternVL3 78B 172.9 91.7 85.0 96.3 82.1 92.9 83.3 66.7 80.0 80.0 76.7 66.7
[Uncaptioned image]  Qwen2.5-VL 3B 113.7 85.0 85.0 66.3 72.3 90.1 66.7 43.3 50.0 40.0 53.3 26.7
[Uncaptioned image]  Qwen2.5-VL 7B 148.2 92.5 82.5 96.3 147.5 73.8 91.6 43.3 56.7 63.3 76.7 63.3
[Uncaptioned image]  Qwen2.5-VL 72B 157.6 95.0 86.65 96.3 70.4 79.8 73.3 60.0 63.3 80.0 83.3 50.0
4.2 Main Results

### v2

@@ III-C SEM-based Evaluation Framework for Benchmark
where $n$ denotes the total number of task indicators in the benchmark. This aggregate score $D_{\mathrm{valid}}\in(0,1]$ reflects the overall of indicator independence across the benchmark. Lower values indicate potential redundancy and multicollinearity issues at a global level, while higher values suggest stronger statistical validity of task indicators as formative constructs.
IV Gold Benchmark
IV-A Motivation
SEM analysis within existing benchmarks reveal that many tasks are characterized by low contribution to latent capability measurement and high redundancy with respect to other tasks. This observation raises a central concern: whether the increasing number of tasks truly enhances evaluation validity, or merely inflates benchmark scale without substantial benefit. A natural idea is to remove these redundant and low-contribution tasks. However, this immediately leads to a methodological challenge: would eliminating such tasks compromise the benchmark’s ability to capture and differentiate model capabilities? If so, task reduction would risk undermining benchmark reliability; if not, it could provide a principled way to improve evaluation efficiency.
To address this issue, we first empirically verify that reducing the number of tasks does not harm the benchmark’s ability to represent model capabilities. However, we also observe that redundancy persists across capability dimensions, indicating that task-level pruning alone is insufficient. This highlights the necessity of introducing a theoretically grounded capability stratification. Therefore, we incorporate a hierarchical structure of capabilities inspired by Piaget’s cognitive development theory, which provides a principled foundation for disentangling and organizing evaluation dimensions.
TABLE IV: Refined Task Selection Results on BLINK
Capability Representative Task VIF Outer Loading
Low-level Relative Depth 1.248 0.76
Visual Correspondence 1.248 0.921
Art Style 2.884 0.869
Mid-level Jigsaw 2.730 0.816
Spatial Relation 1.318 0.829
Forensic Detection 2.137 0.888
High-level Semantic Correspondence 1.962 0.881
Visual Similarity 2.015 0.852
IV-B Empirical Evidence
TABLE V: Reliability and Validity Metrics for Latent Constructs
Benchmark Construct Cronbach’s $\alpha$ Reliability Composite Reliability Convergent Validity
Attribute Reasoning 0.791 0.806 0.905 0.826
Coarse Perception 0.884 0.898 0.945 0.895
Fine-grained perception 0.841 0.945 0.924 0.858
MMBench Cross-instance Perception 0.746 0.940 0.879 0.785
Logical Reasoning 0.890 0.914 0.947 0.900
Relation Reasoning 0.843 0.858 0.927 0.864
high-level 0.846 0.856 0.906 0.763
BLINK low-level 0.717 0.726 0.831 0.713
mid-level 0.800 0.838 0.876 0.703
Experimental Setup. We refine both MMBench [38] and BLINK [34] by removing tasks with $\mathrm{VIF}>5$ or $\mathrm{outerloadings}<0.8$ [146]. Here, outer loading refers to the correlation between an observed indicator (task) and its assigned latent construct, with higher values indicating that the task reliably reflects the underlying capability dimension. This ensures that only non-redundant and sufficiently representative tasks are retained for further SEM-based evaluation. In addition, we ensure that each latent capability dimension retains at least two tasks, which is necessary for computing VIF and outer loadings. If all tasks within a dimension fail to meet the VIF or outer loading thresholds, we retain the tasks that are relatively closest to meeting these criteria, so that each capability dimension remains represented in the SEM analysis.
Task Refinement Results. Tables III and IV summarize the retained tasks with their corresponding VIF and outer loadings. After refinement, all tasks achieve acceptable VIF values and high outer loadings. Especially, the VIF of the Physical Relation task decreases from $2.30$ to $1.75$ while its outer loading increases from $0.872$ to $0.924$, and similar improvements are observed across other tasks, achieving a better balance between multicollinearity and task representativeness. These results demonstrate that SEM filtering successfully removes redundant and low-contribution tasks while maintaining reliable indicators for latent capability measurement.
Reliability and Validity Results. Table V reports the reliability and validity metrics for the latent constructs across MMBench and BLINK. Most constructs achieve acceptable thresholds for Cronbach’s $\alpha$ (>0.70>0.70[147], Composite Reliability (>0.80>0.80[148], and Convergent validity (>0.70>0.70[149], confirming that the retained tasks provide consistent and reliable measurements of the intended capabilities. Formally, Cronbach’s $\alpha$ is defined as:
$\alpha=\frac{k}{k-1}\left(1-\frac{\sum_{i=1}^{k}\sigma_{i}^{2}}{\sigma_{T}^{2}}\right),$ (6)
where $k$ is the number of items, $\sigma_{i}^{2}$ is the variance of item $i$, and $\sigma_{T}^{2}$ is the variance of the total score.
Composite reliability (CR) is calculated as:
$\mathrm{CR}=\frac{(\sum_{i=1}^{k}\lambda_{i})^{2}}{(\sum_{i=1}^{k}\lambda_{i})^{2}+\sum_{i=1}^{k}\theta_{i}},$ (7)
where $\lambda_{i}$ denotes the standardized factor loading of item $i$, and $\theta_{i}$ is its error variance.
Convergent validity, typically assessed by the Average Variance Extracted (AVE), is given by:
$\mathrm{AVE}_{j}=\frac{\sum_{k}\left[w_{p_{i}}^{2}\mathrm{var}\left(\xi_{j}\right)\right]}{\sum_{k}\left[w_{p_{i}}^{2}\mathrm{var}\left(\xi_{j}\right)\right]+\sum_{k}\left(1-w_{p_{i}}^{2}\right)},$ (8)
where $w_{p_{i}}$ denotes the standardized factor loading of the $i$-th indicator on latent construct $j$, and $\mathrm{var}(\xi_{j})$ represents the variance of the latent construct.
For MMBench, constructs such as Logical Reasoning ($\alpha=0.890$, composite reliability =0.947=0.947, convergent validity =0.900=0.900) demonstrate strong internal consistency and convergent validity, indicating well-defined and coherent capability dimensions. For BLINK, while high-level capabilities exhibit good reliability and validity, the low-level dimension shows relatively weaker values ($\alpha=0.717$, convergent validity =0.713=0.713), suggesting room for further refinement. The refined benchmarks capture capability dimensions with satisfactory reliability and validity, making them suitable for systematic evaluation without redundancy.
Figure 5: HTMT heatmap of MMBench latent capabilities.
TABLE VI: HTMT Values of BLINK Latent Capabilities
BLINK High-level Low-level Mid-level
BLINK 1.00 0.50 0.68 0.36
High-level 0.50 1.00 1.12 0.98
Low-level 0.68 1.12 1.00 1.08
Mid-level 0.36 0.98 1.08 1.00
Figure 6: Rank correlation between original benchmarks (blue), SEM-refined benchmarks (red), and human evaluation (green). The four plots correspond to: (i) MMBench Top-50 models, (ii) MMBench Middle-50 models, (iii) BLINK Top-50 models, and (iv) BLINK Middle-50 models.
Discriminant Validity. At the same time, an analysis of HTMT scores reveals that discriminant validity across latent capabilities remains limited. For MMBench as shown in Fig. 5, the cross-capability correlation between Cross-instance Perception and Attribute Reasoning increases from $1.03$ in the original benchmark to $1.20$ after SEM-based refinement, well above the recommended threshold of $0.90$. Similarly, as shown in Table VI BLINK retains high HTMT values, confirming that cross-level dependencies persist. This increase indicates that task-level pruning did not reduce inter-capability entanglement; rather, the latent capability dimensions have become even more overlapping. Such persistent and elevated correlations highlight the limitations of task-level refinement alone and underscore the need for a theoretically grounded capability hierarchy to improve discriminant validity.
TABLE VII: Spearman Rank Correlations Between Original, SEM-refined of MMBench/BLINK and Human Evaluation
Benchmark Subset Origin vs. SEM Origin vs. Human SEM vs. Human
MMBench Top-50 0.9099 0.4193 0.4975
Bottom-50 0.9314 0.2766 0.3376
BLINK Top-50 0.8395 0.4335 0.6138
Bottom-50 0.8508 0.1826 0.2401
Ranking Stability. Ranking stability analysis is presented in Fig. 6, which illustrates the consistency of model rankings before and after task pruning across different benchmarks. As shown in Table VII, for MMBench, the correlation between original and SEM-refined benchmarks is $0.91$ for the top-50 models and $0.93$ for the bottom-50 models, indicating that task refinement preserves the overall ranking structure. The correlations with human evaluation increase from $0.42$ to $0.50$ for the top-50 models and from $0.28$ to $0.34$ for the bottom-50 models, showing improved alignment with human judgment. Similarly, for BLINK, the original vs. SEM-refined correlations are $0.84$ (top-50) and $0.85$ (bottom-50), while correlations with human evaluation rise from $0.43$ to $0.61$ for the top-50 models and from $0.18$ to $0.24$ for the bottom-50 models. Overall, these results demonstrate that SEM-based task filtering maintains ranking stability while increasing the benchmark’s consistency with human evaluation, validating both the effectiveness and reliability of the refined benchmark.
IV-C Hierarchical Structure Based on Piaget’s Developmental Theory
Jean Piaget’s theory of cognitive development is a seminal framework in developmental psychology, which describes how human cognitive abilities evolve through structured stages from infancy to adulthood. The classic stages include the Sensorimotor Stage (0–2 years), characterized by sensory and motor interactions with the environment; the Preoperational Stage (2–7 years), marked by symbolic thinking and basic memory development; the Concrete Operational Stage (7–11 years), in which logical reasoning about concrete events emerges; and the Formal Operational Stage (11+ years), defined by abstract and hypothetical reasoning.
Recent studies [150, 151] suggest that MLLMs exhibit cognitive capabilities roughly equivalent to the Concrete Operational stages of Piaget’s hierarchy. This indicates that MLLMs primarily demonstrate perceptual processing, memory encoding and retrieval, and elementary reasoning (e.g., object classification, spatial relations, and cause-effect inference), while higher-order abstract reasoning and metacognition associated with the Formal Operational Stage remain largely unattainable. Thus, Piaget’s framework provides a natural, theory-driven lens for organizing and evaluating MLLM capabilities, as it aligns with the current developmental scope of these models and offers interpretable cognitive layers for benchmarking.
@@ IV-C Hierarchical Structure Based on Piaget’s Developmental Theory
Motivated by this observation, we propose a new capability hierarchy grounded in Piaget’s theory of cognitive development. Unlike previous heuristic-based groupings, the proposed approach introduces a human-centered, theory-driven structure to model the cognitive underpinnings of MLLM performance. By anchoring the evaluation framework in developmental principles, we aim to achieve a more interpretable, hierarchical, and cognitively faithful representation of model abilities.
@@
This Piaget-inspired cognitive structure is embedded into our benchmark design, with each task aligned to one of the three hierarchical cognitive layers (Perception, Memory, and Reasoning). To ensure that this theoretically grounded stratification is not only conceptually meaningful but also empirically valid, we employ SEM to assess its reliability and distinctiveness. SEM confirms the effectiveness of this organization, demonstrating improved dimensional separability, reduced task redundancy, and stronger measurement consistency compared to conventional heuristic groupings.
IV-D Gold Construction
To construct a human-aligned benchmark structure, we map tasks from existing multimodal datasets to cognitive processes defined in a Piaget-inspired framework (Section IV-C). This process is formalized by using SEM, as outlined in Algorithm 1. We begin by curating an indicator pool by collecting tasks from four widely adopted multimodal benchmarks: MME [152], MMMU-Bench [88], BLINK [34] and MMBench [38]. Each task is manually annotated and assigned to one cognitive layer based on its processing demands (e.g., OCR $\rightarrow$ Perception; landmark recall $\rightarrow$ Memory; domain‐specific inference $\rightarrow$ Reasoning).
We then apply PLS-SEM to estimate both the measurement and structural models in a unified, variance-based manner. Unlike covariance-based SEM, which emphasizes global model fit, PLS-SEM prioritizes maximizing explained variance of latent constructs and predictive relevance of the indicators. The algorithm proceeds iteratively through the following steps.
For each construct $\xi_{j}$, a latent score at iteration $t$ is obtained as a weighted composite of its observed indicators ${x_{ij}}{i=1}^{p_{j}}$:
$\hat{\xi}_{j}^{(t)}=\sum_{i=1}^{p_{j}}\lambda_{ij}^{(t)}x_{ij},\quad\text{with }\sum_{i=1}^{p_{j}}(\lambda_{ij}^{(t)})^{2}=1.$ (9)
Here, initial weights are set uniformly, and are subsequently refined based on inner and outer approximations.
The inner approximation aggregates information from adjacent constructs in the structural model. Specifically, a proxy $z_{j}^{(t)}$ is constructed as
$z_{j}^{(t)}=\sum_{k\in\mathrm{Neigh}(j)}\alpha_{jk}^{(t)}\hat{\xi}k^{(t)},$ (10)
where $\alpha{jk}^{(t)}=\mathrm{sgn}\big(\mathrm{corr}(\hat{\xi}_{j}^{(t)},\hat{\xi}_{k}^{(t)})\big)$ preserves the empirical directionality of associations.
The outer approximation updates the weights through an Ordinary Least Squares regression of indicators on the proxy score:
$\lambda_{ij}^{(t+1)}=\frac{\mathrm{cov}(x_{ij},z_{j}^{(t)})}{\sqrt{\mathrm{var}(x_{ij}),\mathrm{var}(z_{j}^{(t)})}}.$ (11)
The weights are then normalized to unit length for numerical stability:
$\lambda_{ij}^{(t+1)}\leftarrow\frac{\lambda_{ij}^{(t+1)}}{\sqrt{\sum_{k=1}^{p_{j}}(\lambda_{kj}^{(t+1)})^{2}}}.$ (12)
The procedure alternates between latent score estimation and weight updating until convergence, defined as
$\max_{i,j},|w_{ij}^{(t+1)}-w_{ij}^{(t)}|<\epsilon,$ (13)
or equivalently, stability in both indicator weights and structural path coefficients.
Upon convergence, the outer loadings
$\lambda_{ij}=\frac{\mathrm{cov}(x_{ij},\hat{\xi}_{j})}{\sqrt{\mathrm{var}(x_{ij}),\mathrm{var}(\hat{\eta}_{j})}}$ (14)
quantify the reliability of each indicator, while structural path coefficients are obtained by regressing each endogenous construct on its predictors via OLS. These coefficients capture the hypothesized causal relations among latent capabilities, whereas the loadings inform indicator validity and redundancy.
Data: VIF threshold $\delta_{\text{VIF}}$, outer loading threshold $\lambda_{\min}$
1ex// Stage 1: Initialization
Define latent constructs: Perception, Memory, Reasoning;
Assign each task $x_{i}\in\mathcal{X}$ to its latent construct;
Initialize weights $w\leftarrow 1$;
1ex// Stage 2: Iterative Refinement
while task with $\mathrm{VIF}>\delta_{\text{VIF}}$ or $\mathrm{outer\ loading}<\lambda_{\min}$ do
   1. Estimate measurement model via PLS-SEM (maximize explained variance);
   2. Remove redundant tasks with $\mathrm{VIF}>\delta_{\text{VIF}}$;
   3. Remove low-quality tasks with $\mathrm{outer\ loading}<\lambda_{\min}$;
1ex// Stage 3: Validation
1. Validate model reliability using Cronbach’s $\alpha$;
2. Assess discriminant validity using HTMT;
Output: Theory-aligned benchmark structure with validated task indicators
@@ 2. Assess discriminant validity using HTMT;
Figure 7: Structural analysis of Gold by using SEM-based framework.
TABLE VIII: Reliability and Validity Metrics for Latent Constructs
Metrics
Latent Constructs
Memory Perception Reasoning
Cronbach’s $\alpha$ 0.702 0.804 0.902
Reliability 0.732 0.813 0.906
Composite reliability 0.823 0.885 0.927
Convergent validity 0.707 0.720 0.719
SRMR 0.087
$R^{2}$ 0.557
To refine the model, we iteratively prune the indicator pool by removing tasks exhibiting high multicollinearity (variance inflation factor $\delta_{\text{VIF}}$) or insufficient indicator quality (outer loading $\lambda_{\min}$). Following prior empirical standards [153, 154], we set $\delta_{\text{VIF}}=5$ and $\lambda_{\min}=0.75$. This process enhances both statistical validity and conceptual coherence. We further evaluate construct reliability and discriminant validity using Cronbach’s $\alpha$ and the heterotrait–monotrait (HTMT) ratio.
@@
TABLE IX: Comparison of State-of-the-Art Methods on Gold
Methods
Model
size
Overall Perception Memory Reasoning
@@ Methods
[Uncaptioned image]  GPT-4o - 216.3 92.5 92.5 96.3 72.6 91.0 73.3 56.7 73.3 80.0 76.7 70.0
[Uncaptioned image]  Gemini-2.0-Pro - 212.9 90.0 82.5 88.8 81.0 92.8 83.3 60.0 70.0 83.3 83.3 56.7
[Uncaptioned image]  Gemini-2.0-Flash - 215.5 95.0 82.5 96.3 73.5 84.9 60.0 60.0 83.3 80.0 86.7 60.0
[Uncaptioned image]  Gemini1.5-Pro - 183.9 90.0 77.5 68.8 72.8 74.6 63.3 46.7 66.7 70.0 53.3 63.3
@@
[Uncaptioned image] DeepSeek-VL2-tiny 3B 156.2 66.7 71.7 81.3 63.5 84.3 40.0 40.0 40.0 53.3 53.3 50.0
[Uncaptioned image] DeepSeek-VL2-small 16B 165.8 82.5 70.9 85.0 69.0 84.9 53.3 53.3 53.3 36.7 46.7 53.3
[Uncaptioned image] DeepSeek-VL2 27B 173.8 90.0 70.9 73.8 64.9 86.1 66.7 60.0 60.0 36.7 53.3 56.7
[Uncaptioned image] DeepSeek 1B 118.2 62.5 56.7 40.0 62.3 78.8 46.7 30.7 46.7 36.7 30.0 16.7
[Uncaptioned image] DeepSeek 7B 140.7 85.0 79.2 47.5 58.6 82.3 56.7 27.3 43.3 36.7 43.3 33.3
[Uncaptioned image]  Gemma3 4B 148.4 54.2 55.0 81.3 62.5 80.1 53.3 36.7 46.7 46.7 53.3 50.0
[Uncaptioned image]  Gemma3 12B 180.4 75.0 75.0 85.0 75.6 88.3 63.3 33.3 60.0 73.3 63.3 53.3
[Uncaptioned image]  Gemma3 27B 194.2 73.4 72.5 92.5 64.4 86.5 70.0 40.0 70.0 70.0 86.7 63.3
[Uncaptioned image]  InternVL3 2B 161.8 86.7 77.5 85.0 78.1 84.5 30.0 40.0 40.0 56.7 50.0 40.0
[Uncaptioned image]  InternVL3 9B 201.6 89.2 86.7 96.3 80.9 87.9 63.3 66.7 63.3 60.0 73.3 56.7
[Uncaptioned image]  InternVL3 38B 209.9 86.7 82.5 96.3 82.4 92.3 76.7 50.0 70.0 76.7 83.3 63.3
[Uncaptioned image]  InternVL3 78B 221.1 91.7 85.0 96.3 82.1 92.9 83.3 66.7 80.0 80.0 76.7 66.7
[Uncaptioned image]  KimiVL 16B 183.9 92.5 85.0 77.5 72.8 89.5 53.3 46.7 63.3 73.3 63.3 53.3
[Uncaptioned image]  Qwen2.5-VL 3B 160.1 85.0 85.0 66.3 72.3 90.1 66.7 43.3 50.0 40.0 53.3 26.7
[Uncaptioned image]  Qwen2.5-VL 7B 195.9 92.5 82.5 96.3 73.8 91.6 63.3 43.3 56.7 63.3 76.7 63.3
[Uncaptioned image]  Qwen2.5-VL 72B 207.2 95.0 86.7 96.3 70.4 79.8 73.3 60.0 63.3 80.0 83.3 50.0
IV-E Statistics.
The final Gold Benchmark contains 12 tasks, selected from an initial pool of 85 across the four benchmarks. Our analysis follows the guidelines proposed by J. Hair [155]. For measurement model, we examine the indicator weights to evaluate their contribution to the latent constructs. As depicted in the Fig. 7, the weights are consistently greater than 0.75. This outcome indicates that the retained indicators possess strong explanatory power and are effectively capturing the underlying constructs they are intended to measure.
We also assess multicollinearity using VIF. As shown in Fig. 7(b), all indicators yield VIF values well below the conservative threshold of 5 (ranging from 1.12 to 2.75), suggesting the absence of problematic multicollinearity and supporting the discriminant structure of the model. Finally, we evaluate discriminant validity using the Heterotrait-Monotrait (HTMT) ratio in Fig. 7(c). All pairwise HTMT values are below the recommended cutoff of 0.90, with the highest observed value being 0.86 (between Memory and Perception). This result confirms that each latent construct captures a distinct cognitive domain and is not confounded with others.
To further verify the statistical quality of the benchmark, we conduct standard reliability and validity tests on the latent constructs. As shown in Table VIII, all constructs meet the recommended thresholds for internal consistency and convergent validity. Specifically, Cronbach’s $\alpha$ values are all above 0.70, reflecting acceptable internal reliability. Composite reliability scores exceed 0.80 across all dimensions, confirming the coherence of the grouped indicators. In terms of convergent validity, all constructs exhibit Average Variance Extracted (AVE) values above 0.60, suggesting that each construct explains a substantial portion of the variance in its associated indicators.
In addition to reliability and validity metrics, we also evaluate the standardized root mean square residual (SRMR) and the $R^{2}$ value for each latent construct, as presented in Table VIII. SRMR provides an absolute measure of model fit, indicating how well the observed data align with the theoretical model. For the Perception construct, SRMR is 0.087, slightly exceeding the conventional threshold of 0.08, suggesting an overall good fit. The $R^{2}$ value reflects the explanatory power of the model, with Perception achieving $R^{2}=0.557$, indicating that more than half of the variance in its indicators is accounted for by the latent construct. Memory and Reasoning also exhibit strong explanatory power, with Cronbach’s $\alpha$, composite reliability, and convergent validity all exceeding recommended thresholds. Together, these metrics confirm that the Gold benchmark possesses robust reliability, valid measurement of latent constructs, and sufficient explanatory power to capture the underlying cognitive dimensions.
V Experimental Study
V-A Experiment Setting
To evaluate the quality and interpretability of the Gold benchmark, we design a series of experiments involving both open-source and closed-source MLLMs. Our evaluation includes two core settings: (1) Main Experiment, where we assess mainstream models using a standardized pipeline to reveal overall performance differences; (2) SEM-based Benchmark Evaluation Experiment, where we validate Gold’s diagnostic effectiveness by measuring key metrics such as dimensional diversity ($D_{\mathrm{div}}$), task contribution ($\mathrm{TC}$), and indicator validity ($D_{\mathrm{valid}}$). We further assess its interpretability by comparing these outputs with human ratings to examine alignment with human judgment.
Evaluation Models. For main experiment, we evaluate both closed-source and open-source MLLMs. For the closed-source group, we include GPT-4o [4], Gemini-2.0 [156] and Gemini-1.5 [157]. For the open-source group, we consider a range of models with varying scales, including DeepSeekVL2 [158], DeepSeek [159], Gemma3 [160], InternVL3 [132], KimiVL [161] and Qwen2.5-VL [131]. These models represent the current frontier of multimodal LLM development and cover diverse architecture families and parameter sizes. For the benchmark evaluation experiment, we leverage evaluation results from 165 multimodal language models collected via VLMEvalKit [142], which provides standardized testing across multiple benchmarks. The Metrics used in the experiment will be described in detail in the Appendixes.
V-B Main Results